OpenStudy (dr0zier99):

What is the equation of a line passing through (-3, 7) and having a slope of - 1/5 y = -5x + 22 y = 5x + 22

pooja195 (pooja195):

First of all, remember what the equation of a line is: y = mx+b Where: m is the slope, and b is the y-intercept To start, you know what m is; it's just the slope, which you said was -1/5. So you can right away fill in the equation for a line somewhat to read: y=-1/5x+b.

pooja195 (pooja195):

To find the y intercept (b) We need to plug in the points \[\huge~\rm~ 7=-1/5x-3+b\] isolate b

pooja195 (pooja195):

Or you can solve for b like this: b=7-(-1/5)(-3)

OpenStudy (dr0zier99):